Minh and Mikayla are selling wrapping paper for a school fund-raiser. Customers can buy either packages of folded paper or rolls of paper. Minh sold 6 rolls of paper and 2 packages of folded paper for a total of $44. Mikayla sold 3 rolls of paper and 9 packages of folded paper for a total of $54. How much does each roll cost? How mush does each package of folded paper cost?
Answer by jorel555(1290)

Let x=rolls of paper,y=folded paper:
6x+2y=44
3x+9y=54
-------
6x+2y=44
6x+18y=108
16y=108-44=64
y=4
x=6
Thus, a roll of paper cost $6,folded paper was $4 each!!!!!!!